OpenStudy (anonymous):

Solve the inequality. 4 + |t + 2| < 11 A. t < 5 and t > 9 B. t > 5 or t < –9 C. –5 < t < 9 D. –9 < t < 5

OpenStudy (anonymous):

\[\left| t+2 \right|<11-4,.\left| t+2 \right|<7\] \[\left| x+a \right|<b,-b<x+a<b\] subtract a from each.

OpenStudy (anonymous):